A firm faces a 30 percent tax rate and has $500m in assets, currently financed entirely with equity. Equity is worth $100 per share, and book value of equity is equal to market value of equity. Also, let's assume that the firm's expected EBIT is $60m. The firm is considering switching to a 25 percent debt capital structure, and has determined that they would have to pay a 10 percent yield on perpetual debt. How much will ROE change if they switch to the proposed capital structure?

Top of Form

Multiple Choice

  • There will be no change in the firm's ROE.
  • The ROE will increase by 0.47 percent.
  • The ROE will increase by 1.15 percent.

The ROE will increase by 0.82 percent.


1
Expert's answer
2021-06-18T06:27:38-0400

Interest expense"= \$500M X 0.25 X 0.10"

"=\$12.5M" per year

Net income"= [\$60M - \$12.5M] X (1 - 0.3)"

"= \$47.5M X 0.7"

"= \$33.25M"

Shareholder's equity "= \$500M - (25\% \space of \space \$500M)"

"= \$375M"

New return on equity "=\frac{ Net \space income}{ Shareholder's\space equity}"

"= \frac{\$33.25M}{ \$375M}"

"= 0.088666666 or 8.8666666\%"

Old return on equity"=\frac{\$60M \times (1 - 0.3) }{ \$500M}"

"= \frac{\$49M }{\$500M}"

"= 0.084\space or \space8.4\%"

Change in return on equity = New return on equity - Old return on equity

"= 8.8666666\% - 8.4\%"

"= 0.4666666\%"

The return on equity will increase by"\space 0.47\%."